# Has your pretrained model improved? a multi-head posterior based approach

### 02 Jan 2024 v1

@@ Abstract
The emergence of pretrained models has significantly impacted from Natural Language Processing (NLP) and Computer Vision to relational datasets. Traditionally, these models are assessed through fine-tuned downstream tasks. However, this raises the question of how to evaluate these models more efficiently and more effectively. In this study, we explore a novel approach where we leverage the meta features associated with each entity as a source of worldly knowledge and employ entity representations from the models. We propose using the consistency between these representations and the meta features as a metric for evaluating pretrained models. Our method’s effectiveness is demonstrated across various domains, including models with relational datasets, large language models and images models.
@@ 1 Introduction
Pre-training on large models is becoming increasingly common in various machine learning applications, thanks to the growing amount of user-generated content. This is evident in areas such as Natural Language Processing (NLP) with models like GPT (Generative Pretrained Transformer), and in the vision-language domain with models like CLIP. Typically, the effectiveness of these models is evaluated using downstream tasks. However, these can be relatively costly if all tasks need to be performed. Is there a more efficient, yet simple method to evaluate these models?
Entity representations as known as embeddings, generated from these models, can be utilized directly or indirectly by downstream tasks and can also be fine-tuned as needed. The associated meta features with these embeddings can be considered as the model’s foundational knowledge of the world it’s learning from. This could be the class category for image data or semantic and syntactic information for words. Despite having the same meta features, embeddings differ across models. The degree of consistency between the embeddings and meta features can therefore serve as a performance metric for model evaluation.
The embedding space is complex and challenging to interpret or explain. Despite extensive efforts to decipher it, the intricacies of the embedding space go beyond mere linear interpretability as some research suggests. In this research, we hypothesize that the embeddings reside within a manifold space where Euclidean distance is not an appropriate metric for gauging the similarity between two embeddings. Meta features are capable of grouping these embeddings into clusters, we assume each forming a sub-manifold space. When the clusters are fine granulated enough, we can assumingly use a Gaussian distribution to approximate each cluster. Taken together, these clusters constitute a mixture of Gaussian distributions. By calculating the posterior probabilities of the entities, we can calculate the consistency of the meta features and embeddings.
In this study, we introduce a unique approach to evaluate the performance of pretrained models. Specifically, we:
@@
    Adopt a distinct perspective towards the model. Instead of focusing solely on downstream performance, we place greater emphasis on the quality of the entities’ representations that the model can generate.
@@
    Consider the features associated with the entity representations as the benchmark to assess their quality. We hypothesize that the meta features can partition the embedding space into distinct clusters. The quality of these clusters can be evaluated using the posterior probability of Gaussian mixture models.
@@
    While there are multiple methods to interpret this meta feature spaces, we present a tree-based approach as an example that uses meta features to segment entities into clusters.
@@
    Test the effectiveness of our proposed method on a variety of datasets across different domains, ranging from recommendation-based models to language models and image models. We present both qualitative and quantitative evidence to demonstrate the approach’s efficacy.
@@ 2 Algorithm Framework
For a given domain, we first collect a large size of entities with rich meta features. Then for any given pretrained model, we can generate an embedding dataset denoted as $\mathbf{X}={\mathbf{x}_{1},...,\mathbf{x}_{N}}$, where each $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and $1\leq i\leq N$. Here, $N$ represents the number of entities and $d$ signifies the dimension of the embeddings. Simultaneously, we can produce a corresponding feature set $\mathbf{F}={\mathbf{f}_{1},...,\mathbf{f}_{N}}$. Each feature $\mathbf{f}_{i}$ comprises both categorical and numerical features. We convert numerical features into categorical ones for consistency. The primary objective is to examine the consistency between these two datasets, $\mathbf{X}$ and $\mathbf{F}$.
In the simplest scenario where $\mathbf{f}_{i}$ has only one feature, a straightforward method of segmentation is to form clusters based solely on these features, with each category creating its own cluster. However, when $\mathbf{f}_{i}$ has $m$ features, and each feature has $k_{j}$ categories, the number of combinations becomes $\prod_{j=1}^{m}k_{j}$. This results in a significantly larger number of clusters. We will postpone the discussion on finding the best split to a later session. In the upcoming discussion, we will assume that we already have a split criterion for the dataset $\mathbf{X}$.
2.1 Posterior Based Embedding Evaluating Metric
We aim to evaluate the effectiveness of different models that generate these embeddings to determine the best one. The splitting criteria $\mathbf{S}$ can divide the entities into a group of clusters $C_{1},C_{2},...,C_{n}$, with each entity belonging to a single cluster, where $n$ is number of clusters. To evaluate the quality of the cluster, we adopt a posterior based method.
In the context of GMM, it is assumed that the data is generated from a combination of multiple Gaussian distributions. Each component of this mixture corresponds to one of the distinct clusters within the dataset. For any given data point within a particular cluster, the formula for calculating its posterior probability in the GMM framework can be expressed as follows:
$P(\theta=k|\mathbf{x}_{k})=\frac{P(\mathbf{x}_{k}|\theta=k)P(\theta=k)}{\sum_{j=1}^{m}P(\mathbf{x}_{k}|\theta=j)P(\theta=j)},$ (1)
where $\mathbf{x}_{k}$ represents every point in cluster $k$. $P(\theta=k)$ = number of pionts in cluster $k$ / $N$.
To assess the quality of embeddings $\mathbf{X}$ within the context of a splitting $\mathbf{S}$, we compute the overall evaluation metric $P_{\mathbf{S}}^{\mathbf{X}}$ by averaging the log probabilities of all the embeddings across all clusters. This metric provides an assessment of the quality of the embeddings. We call it the average of log posterior.
@@ 2.1 Posterior Based Embedding Evaluating Metric
One challenge with the aforementioned formula is its high sensitivity to outliers. A single outlier could lead to an extremely large value for $P_{\mathbf{S}}^{\mathbf{X}}$. To mitigate the impact of such outlier entities, we implement a clipping mechanism for embeddings with very small posterior probabilities. Specifically, if $P(\theta=k|\mathbf{x}_{k})$ is less than $k/N*\varepsilon$, we exclude the entity from the $P_{\mathbf{S}}^{\mathbf{X}}$ computation.
Another challenge arises when the embeddings exist in large dimensions. If the number of embeddings in each cluster is smaller than the embedding dimension, this leads to a rank-deficient covariance. To address this, we propose a multi-head solution. In this approach, we randomly select $v$ dimensions and evaluate based on these dimensions. This process is repeated multiple times and the average results are used. This concept is inspired by the random forest algorithm (Breiman, 2001). Additionally, we apply the routine regularization approach, i.e., adding $\epsilon$$I$ to the covariance matrix. The value of $\epsilon$ is decided in the following manner.
@@ 2.2 One Meta Feature Based Clustering
In the simplest scenario, where the feature vector $\mathbf{f}_{i}$ consists of only one feature, a clear and intuitive approach to segmentation is to form clusters based solely on these features. This method is straightforward as it capitalizes on the inherent characteristics of the data. Each unique category within the data forms its own distinct cluster, effectively grouping similar items together. The consistency of these clusters with the embeddings can serve as a measure of the quality of the embeddings. However, it’s important to note that extending this approach to accommodate more than two meta features is not as straightforward.
@@ 2.3 Meta features + representation based segmentation
Inspired by EmbeddingTree algorithm, we can construct the a tree based on the entities and all the leaf nodes are the the finally clusters, specifically: we first convert non-binary categorical features into binary ones by asking yes-no questions regarding to each of their categorical values and get the binary feature sets: $\mathbf{G}=\{\mathbf{g}_{1},...,\mathbf{g}_{N}\}$ ($\mathbf{g}_{i}\in\{0,1\}^{q}$, $1\leq i\leq N$), $q$ denote the total number of converted binary features.
With the processed data, we describe the EmbeddingTree algorithm with details in Algorithm 1. We iterate through the $q$ features (line 6) and evaluate them based on the splitting criteria described in Section 2.4 to pick out the best feature for splitting (line 8-10), using the feature’s binary value (line 11-13). The above procedure is executed recursively (line 15-16) until the splitting criterion ($\Theta$), e.g., the number of entities per tree node or the tree depth, is no longer satisfied (line 2). With the given embedding and feature data, the whole procedure is deterministic.
@@
One critical component of Algorithm 1 is the criterion used to select the best splitting feature. The criterion is computed based on the approximate MAP for GMMs inspired by (Zheng et al., 2023).
We assume the embedding can be modeled as two mixture Gaussians. The expectation-maximization (EM) algorithm is used to jointly estimate all the parameters and latent variables. The latent variables, $z_{i,j}$, denote the probability that sample $i$ is in cluster $j$. With $N$ as the number of observations and $J$ be the number of Gaussian clusters (in this case, $J=2$), $z=\{z_{1,1},z_{1,2},...,z_{N,J-1},z_{N,J}\}$, the complete likelihood (including the latent variables) is:
@@
To find the best binary feature that splits the embedding and forms the best GMM, we go through every feature. Each candidate binary feature splits the embeddings into two clusters, each is then formulated as a Gaussian. For each feature, suppose the first $s$ embeddings have feature value $F^{k}=0$ and the rest $N-s$ embeddings have feature value $F^{k}=1$. We estimate the weights, means, and variances for both clusters using the maximum likelihood estimation (MLE).
@@
In other words, rather than the soft clustering of GMM, our algorithm performs a hard clustering. Thus, if $\mathbf{x}_{i}$ is in cluster $j$, then $z_{i,j}=1$ and $z_{i,j^{\prime}}=0$ for all $j\neq j^{\prime}$. Given this approximation, the likelihood can be obtained by summing over the $z$:
@@
We can treat each split feature as another random variable $\theta$, to choose the best split feature, we want to maximum $P(\mathbf{x},\mu,\Sigma,w,\theta)$, in another words, we want to find $\theta$ gives the largest $P(\mathbf{x},\mu,\Sigma,w)$.
@@ 2.5 Finding the Best Splitting Point
For simplicity, we only consider $\theta$ as the random variable we want to estimate, by injecting the prior information into the formula, we can treat each splitting feature with different weight. By applying Maximum A Posteriori Estimation (MAP), we can formulate the problem as following:
@@
where $q$ is number of possible splits.
@@
By applying this formula, we can apply the prior knowledge on the importance of the feature to find the split that maximize $\log\hat{P}$.
@@ 2.6 Embedding Comparison based on the same splitting criteria
3 Related Work
3.1 Pretrained models
Vision-Language Pre-training (VLP): With the rapid expansion of model capacity and computational resources, the input to deep neural networks has evolved beyond single modality such as text or image. Vision-Language Pre-training (VLP) was introduced to bridge the gap between different modalities, effectively harnessing cross-modality information from both images and text. Leveraging the successful pre-training and fine-tuning paradigm prevalent in NLP, VLP models have demonstrated exceptional performance in complex vision-language tasks. These tasks include image captioning, visual question answering, and visual reasoning. Among the existing studies, a noteworthy contribution is the CLIP (Radford et al., 2021) model, which employs the concept of contrastive learning to align images and text. CLIP simultaneously trains an image encoder and a text encoder on millions of image-text pairs collected from the internet. The resulting encoders have demonstrated impressive performance on downstream tasks due to their zero-shot classification capability.
Pretrained Dual-Transformers(PDT) for Bipartite Graphs: PDT (Dai et al., 2023)focuses on learning contextual knowledge from a user-content interaction dataset, which is depicted as a bipartite graph. The study identifies two key contexts in the graph: user-side and content-side. The goal of learning from these contexts is framed as two contrastive learning tasks and is applied to a recommendation task. Evaluations on two large popular datasets reveal that PDT outperforms baselines in six metrics.
3.2 Embedding Evaluation Metric
Several significant studies have focused on the evaluation of embeddings in Natural Language Processing (NLP). Schnabel et al. Schnabel et al. (2015) conducted a comprehensive evaluation of word embeddings using both intrinsic and extrinsic measures, focusing on specific tasks and downstream applications. Baroni et al. Baroni et al. (2014) compared count-based and predictive models for word representation, using tasks like semantic and syntactic analogies, word similarity, and concept categorization for evaluation. Conneau et al. Conneau et al. (2018) introduced the Cross-lingual Natural Language Inference (XNLI) corpus as a benchmark for multilingual sentence embeddings evaluation. Niven and Kao Niven & Kao (2019) proposed a new evaluation method for embeddings based on their ability to comprehend logical reasoning in natural language. Zhang et al. Zhang et al. (2020) presented BERTScore, an evaluation metric for text generation models based on BERT embeddings that correlates well with human judgment.
Several works have been proposed to evaluate embeddings between different models or various epochs of the same model (Cui et al., 2022; Ding et al., 2021). One such method is GULP (Boix-Adsera et al., 2022), which ensures consistent control over the difference in prediction performance between two representations for regularized linear prediction tasks. GULP exhibits desirable structural properties, including the triangle inequality and invariance under orthogonal transformations. Recently,  (Gurnee & Tegmark, 2023) demonstrated that Language Models (LLMs) have the capacity to learn linear representations of space and time. These representations are consistent across various entities and exhibit a degree of prompt-independence. The study also suggests that individual neurons can exhibit high sensitivity to these features. Furthermore, the research employed out-of-sample $R^{2}$ as a metric for linear probes, substantiating the hypothesis that larger models tend to perform better.

### 14 Feb 2024 v4

# Has your Pre-trained Model Improved? A Multi-head Posterior Based Approach

@@ Abstract
The emergence of pre-trained models has significantly impacted Natural Language Processing (NLP) and Computer Vision to relational datasets. Traditionally, these models are assessed through fine-tuned downstream tasks. However, this raises the question of how to evaluate these models more efficiently and effectively. In this study, we explore a novel approach where we leverage the meta-features associated with each entity as a source of worldly knowledge and employ entity representations from the models. We propose using the consistency between these representations and the meta-features as a metric for evaluating pre-trained models. Our method’s effectiveness is demonstrated across various domains, including models with relational datasets, large language models, and image models.
@@ 1 Introduction
The rise in user-generated content has led to the widespread adoption of pre-training large models across numerous machine-learning fields. This trend is particularly noticeable in Natural Language Processing (NLP) with GPT (Generative Pretrained Transformer) models  (Brown et al., 2020) and vision-language domains with models like CLIP  (Radford et al., 2021). Usually, the performance of these models is assessed through various downstream tasks such as Machine Translation (Wang et al., 2023), Factuality (Chen et al., 2019), Question answering (Liang et al., 2023), Multilingual tasks (Bang et al., 2023; Ahuja et al., 2023), which can be relatively expensive if a large number of tasks are considered necessary for a robust evaluation. Is there an alternative approach that offers both efficiency and simplicity for evaluating these models?
Entity representations, also known as embeddings generated from these models, can be utilized directly or indirectly by downstream tasks and fine-tuned as needed. The associated meta-features with these embeddings can be considered as the model’s foundational knowledge of the world it’s learning from. This could be the class category for image data or semantic and syntactic information for words. Despite having the same meta-features, embeddings differ across models. Therefore, the degree of consistency between the embeddings and meta-features can serve as a performance metric for model evaluation.
Embedding spaces are complex and challenging to interpret or explain. Despite extensive efforts to decipher it, its intricacies go beyond mere linear interpretability, as some research suggests. In this research, we hypothesize that the embeddings reside within a manifold space where Euclidean distance is not an appropriate metric for gauging the similarity between two embeddings. Meta-features are capable of grouping these embeddings into clusters, we assume each forming a sub-manifold space. When the clusters are sufficiently fine-grained, it is possible to approximate each cluster using a Gaussian distribution. Collectively, these clusters form a mixture of Gaussian distributions. By determining the posterior probabilities of the entities, the consistency of meta-features and embeddings can be assessed.
In this study, we introduce a unique approach to evaluate the performance of pre-trained models. Specifically, we:
@@
    Adopt a distinct perspective towards the model. Instead of focusing solely on downstream performance, we emphasize the quality of the entities’ representations that the model can generate.
@@
    Consider the features associated with the entity representations as the benchmark to assess their quality. We hypothesize that the meta-features can partition the embedding space into distinct clusters. The quality of these clusters can be evaluated using the posterior probability of Gaussian mixture models.
@@
    While there are multiple methods to interpret these meta-feature spaces, we present a tree-based approach as an example that uses meta-features to segment entities into clusters.
@@
    Test our proposed method’s effectiveness on various datasets across domains, ranging from recommendation-based to language and image models. We present both qualitative and quantitative evidence to demonstrate the approach’s efficacy.
2 Related Work
This section reviews the literature on three areas: 1) Pre-trained models, 2) Vision-Language Pre-training (VLP), and 3) Pretrained Dual-Transformers (PDT) for Bipartite Graphs.
2.1 Pre-trained models
Vision-Language Pre-training (VLP): With the rapid expansion of model capacity and computational resources, the input to deep neural networks has evolved beyond a single modality, such as text or image. Vision-language pre-training (VLP) was introduced to bridge the gap between different modalities, effectively harnessing cross-modality information from both images and text. Leveraging the successful pre-training and fine-tuning paradigm prevalent in NLP, VLP models have demonstrated exceptional performance in complex vision-language tasks. These tasks include image captioning, visual question answering, and visual reasoning. Among the existing studies, a noteworthy contribution is the CLIP (Radford et al., 2021) model, which employs the concept of contrastive learning to align images and text. CLIP simultaneously trains an image encoder and a text encoder on millions of image-text pairs collected from the internet. The resulting encoders have demonstrated impressive performance on downstream tasks due to their zero-shot classification capability.
Pretrained Dual-Transformers (PDT) for Bipartite Graphs: PDT (Dai et al., 2023) focuses on learning contextual knowledge from a user-content interaction dataset, which is depicted as a bipartite graph. The study identifies two key contexts in the graph: user-side and content-side. The goal of learning from these contexts is framed as two contrastive learning tasks and is applied to a recommendation task. Evaluations of two large popular datasets reveal that PDT outperforms baselines in six metrics.
@@ 3 Algorithm Framework
This section presents the algorithmic framework of our proposed metric for evaluating embeddings.
For a given domain, we first collect a large size of entities with rich meta-features. Then for any given pre-trained model, we can generate an embedding dataset denoted as $\mathbf{X}={\mathbf{x}_{1},...,\mathbf{x}_{N}}$, where each $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and $1\leq i\leq N$. Here, $N$ represents the number of entities, and $d$ signifies the dimension of the embeddings. Simultaneously, we can produce a corresponding feature set $\mathbf{F}={\mathbf{f}_{1},...,\mathbf{f}_{N}}$. Each feature $\mathbf{f}_{i}$ comprises categorical and numerical features. We convert numerical features into categorical ones for consistency. The primary objective is to examine the consistency between these two datasets, $\mathbf{X}$ and $\mathbf{F}$.
In the simplest scenario where $\mathbf{f}_{i}$ has only one feature, a straightforward segmentation method is to form clusters based solely on these features, with each category creating its cluster. However, when $\mathbf{f}_{i}$ has $m$ features, and each feature has $k_{j}$ categories, the number of combinations becomes $\prod_{j=1}^{m}k_{j}$. This results in a significantly larger number of clusters. We will postpone the discussion on finding the best split to a later session. In the upcoming discussion, we will assume that we already have a split criterion for the dataset $\mathbf{X}$.
4 Proposed Method: Posterior Based Embedding Evaluating Metric
We aim to evaluate the effectiveness of different models that generate these embeddings to determine the best one. The splitting criteria, $\mathbf{S}$, can divide the entities into a group of clusters $C_{1},C_{2},...,C_{n}$, with each entity belonging to a single cluster, where $n$ is the number of clusters. To evaluate the quality of the cluster, we adopt a posterior-based method.
In the context of GMM, it is assumed that the data is generated from a combination of multiple Gaussian distributions. Each component of this mixture corresponds to one of the distinct clusters within the dataset. The probability of any given data point, $\mathbf{x}$, belonging to the $k$th cluster is estimated by computing the posterior probability in the GMM framework which can be expressed as follows:
$P(\theta=k|\mathbf{x})=\frac{P(\mathbf{x}|\theta=k)P(\theta=k)}{\sum_{j=1}^{m}P(\mathbf{x}|\theta=j)P(\theta=j)},$ (1)
where $P(\theta=k)$ = number of points in the kth cluster $\frac{\text{number of points in the $k$th cluster }}{N}$
To assess the quality of the embeddings $\mathbf{X}$ within the context of a splitting $\mathbf{S}$, we compute the overall evaluation metric by averaging the log probabilities of all the embeddings across all clusters. This metric provides an assessment of the quality of the embeddings. We call it the average of the log posterior ($ALP$).
@@ 4 Proposed Method: Posterior Based Embedding Evaluating Metric
One challenge with the formula above is its high sensitivity to outliers. A single outlier could lead to an extremely large value for $ALP_{\mathbf{S}}^{\mathbf{X}}$. We implement a clipping mechanism for embeddings with very small posterior probabilities to mitigate the impact of such outlier entities. Specifically, if $P(\theta=k|\mathbf{x}_{k})$ is less than $k/N*\varepsilon$, we exclude the entity from the $ALP_{\mathbf{S}}^{\mathbf{X}}$ computation.
Another challenge arises when the embeddings exist in large dimensions. If the number of embeddings in each cluster is smaller than the embedding dimension, this leads to a rank-deficient covariance. To address this, we propose a multi-head solution. In this approach, each head is a randomly selected $v$ dimensions from the existing $d$ dimensions and $ALP$ is estimated based on these dimensions. This process is repeated multiple times, and the average results are used which we refer to as the Mean of the . This concept is inspired by the random forest algorithm (Breiman, 2001) and Matryoshka Representation Learning (Kusupati et al., 2022). Additionally, we apply the routine regularization approach, i.e., adding $\epsilon$$I$ to the covariance matrix. The value of $\epsilon$ is decided in the following manner.
@@ 4.1 One Meta Feature Based Clustering
In the simplest scenario, where the feature vector $\mathbf{f}_{i}$ consists of only one feature, a straightforward and intuitive approach to segmentation is to form clusters based solely on these features. This method is straightforward as it capitalizes on the inherent characteristics of the data. Each unique category within the data forms its distinct cluster, effectively grouping similar items. The consistency of these clusters with the embeddings can serve as a measure of the quality of the embeddings. However, it’s important to note that extending this approach to accommodate more than two meta-features is not as straightforward.
@@ 4.2 Meta features + representation based segmentation
Inspired by EmbeddingTree algorithm, we can construct the tree based on the entities, and all the leaf nodes are the final clusters, specifically: We first convert non-binary categorical features into binary ones by asking yes-no questions regarding each of their categorical values and get the binary feature sets: $\mathbf{G}=\{\mathbf{g}_{1},...,\mathbf{g}_{N}\}$ ($\mathbf{g}_{i}\in\{0,1\}^{q}$, $1\leq i\leq N$), $q$ denote the total number of converted binary features.
With the processed data, we describe the EmbeddingTree algorithm with details in Algorithm 1. We iterate through the $q$ features (line 6) and evaluate them based on the splitting criteria described in Section 4.3 to pick out the best feature for splitting (line 8-10), using the feature’s binary value (line 11-13). The above procedure is executed recursively (lines 15-16) until the splitting criterion ($\Theta$), e.g., the number of entities per tree node or the tree depth, is no longer satisfied (line 2). With the given embedding and feature data, the whole procedure is deterministic.
@@
One critical component of Algorithm 1 is the criterion for selecting the best splitting feature. The criterion is computed based on the approximate MAP for GMMs inspired by (Zheng et al., 2023).
We assume the embedding can be modeled as two mixture Gaussians. The expectation-maximization (EM) algorithm is used to estimate all the parameters and latent variables jointly. The latent variables, $z_{i,j}$, denote the probability that sample $i$ is in cluster $j$. With $N$ as the number of observations and $J$ as the number of Gaussian clusters (in this case, $J=2$), $z=\{z_{1,1},z_{1,2},...,z_{N,J-1},z_{N,J}\}$, the complete likelihood (including the latent variables) is:
@@
We go through every feature to find the best binary feature that splits the embedding and forms the best GMM. Each candidate binary feature splits the embeddings into two clusters, each formulated as a Gaussian. For each feature, suppose the first $s$ embeddings have feature value $Fk=0$ and the rest $N-s$ embeddings have feature value $Fk=1$. We estimate both clusters’ weights, means, and variances using the maximum likelihood estimation (MLE).
@@
$\hat{w_{2}}=\frac{N-s}{N}.$
In other words, our algorithm performs a hard clustering rather than the soft clustering of GMM. Thus, if $\mathbf{x}_{i}$ is in cluster $j$, then $z_{i,j}=1$ and $z_{i,j^{\prime}}=0$ for all $j\neq j^{\prime}$. Given this approximation, the likelihood can be obtained by summing over the $z$:
@@
We can treat each split feature as another random variable $\theta$. To choose the best-split feature, we want to maximize $P(\mathbf{x},\mu,\Sigma,w,\theta)$; in other words, we want to find $\theta$ that gives the largest $P(\mathbf{x},\mu,\Sigma,w)$.
@@ 4.4 Finding the Best Splitting Point
For simplicity, we only consider $\theta$ as the random variable we want to estimate; by injecting the prior information into the formula, we can treat each splitting feature with a different weight. By applying Maximum A Posteriori Estimation (MAP), we can formulate the problem as follows:
@@
where $q$ is the number of possible splits.
@@
By applying this formula, we can use the prior knowledge of the importance of the feature to find the split that maximizes $\log\hat{P}$.